The Marketing Department reckons that the supply function for Ethyl Alcohol is given by

 Q_x^S=5,000+2P_x-3P_y- P_w

Where Px is the price of ethyl alcohol, Pr represents the price of Isopropyl Alcohol, and Pw is the price of input used to produce ethanol. Given that the amount of ethyl alcohol sells at P40 per unit, isopropyl alcohol is sold for P35 per unit; and the price of inputs is P2,000. How much ethyl alcohol is produced?


1
Expert's answer
2021-01-31T15:12:49-0500
"Q_x^S=5000+2\cdot40-3\cdot35-2000=2975."

2975 units of ethyl alcohol is produced.